Determine the laplace transform

i) ∂2u(x,t)/∂t2

ii)∂2u(x,t)/∂x2

1
Expert's answer
2021-10-22T00:39:26-0400

i)

"L(u(x,t))=\int^{\infin}_0 e^{-st}u(x,t)dt=U(x,s)"


"L(u_t(x,t))=\int^{\infin}_0 e^{-st}u_t(x,t)dt=e^{-st}u(x,t)|^{\infin}_0+s\int^{\infin}_0 e^{-st}u(x,t)dt="

"=sU(x,s)-u(x,0)"


"L(u_{tt}(x,t))=s^2U(x,s)-su(x,0)-u_t(x,0)"


ii)

"L(u_x(x,t))=\int^{\infin}_0 e^{-st}u_x(x,t)dt=U_x(x,s)"


"L(u_{xx}(x,t))=\int^{\infin}_0 e^{-st}u_{xx}(x,t)dt=U_{xx}(x,s)"